1. The amount of time it takes a patient to fully recover from a virus-inflicted illness (in days) follows a normal distribut k=28 n=1 m=15
Answer #1

solutions given dator К= 28 1 - 45 500 төг As As a Qa - slissus ܕܦܠܥܠܪܕܝ fully recoved from the о Х СР, а) <br> to 9ь Aь с ня28 HoM XON CM,45=6) VISHA:M = 56 > 28 Jest statistic (2) z = z - 28 45 1528 where c = 45 2 - 45 - 28 45528 - > 1.999 ,2=0.01